OpenStudy (anonymous):

what is the value of a? -2(a-8)+6a=36 a=?

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OPENSTUDY!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

Since we have parentheses we will need to distribute.

OpenStudy (kidrah69):

\[-2(a-8)+6a=36 \] \[-2\times~a=?\] \[-2\times-8=?\]

OpenStudy (kidrah69):

Any ideas? -.-

OpenStudy (kidrah69):

@kayleebrownell

OpenStudy (anonymous):

sorry don't really know how to work this app, and -2 * -8 = 16 right?

OpenStudy (kidrah69):

yep :) and its ok :P when you do that you end up with.... \[-2a+16+6a=36\] \[6-2=?\]

OpenStudy (kidrah69):

Now we are just combining like terms :)

OpenStudy (anonymous):

6 - 2 = 4

OpenStudy (kidrah69):

\[4a=36\] \[36\div4=?\]

OpenStudy (anonymous):

a = 9?

OpenStudy (kidrah69):

yep :) good job!

OpenStudy (anonymous):

thankyou (:

OpenStudy (kidrah69):

:)

OpenStudy (anonymous):

Remember to close this post when you are done. :)